# A simple, flexible method for timing cross-calibration of space missions

### v1

@@ 3.5 Pipeline
All steps above are automated in the TOAextractor code v. 0.5 (M. Bachetti, 2025b). This code is structured as a pipeline, using the luigi orchestrator [^4] https://luigi.readthedocs.io/en/stable/ v. 3.6.0. This software allows for a simple and modular way to define the steps of the pipeline, and to run them in parallel over multiple files by using multiple “workers”. The end of each step of the pipeline is marked by the production of a file. The appearance of this file triggers the start of the next step. Failing steps do not produce the output file, blocking the pipeline on a given file. Fixing the error and re-running the pipeline will restart the processing from the failed step. To reprocess a given dataset from one particular step of the pipeline, it is sufficient to delete all files produced by the step to be repeated and all subsequent steps, and re-run the pipeline.
3.6 Plotting and quality assurance
@@
Figure 5: Interactive summary plot produced by TOAextractor. This plot allows for a simple exploration of the residuals, by showing diagnostic information as the user hovers over each point. The horizontal grey band indicates the X-ray residual measured by A. H. Rots et al. (2004).Problematic observations and time intervals are easily identified: some observations with large residuals might be due to low count rates, temporary instrumental issues, or local issues with the radio ephemerides (e.g. the large deviation around April 2020, MJDs 58925–58975, due to the closure of the observatory for the COVID pandemic; recent glitches can also make the solution less reliable).
@@
This pipeline is conceived to be run on a large number of observations from different missions, possibly ever increasing over time. As such, it is important to have a quick way to check the quality of the data. For this reason, we developed an interactive summary plot (Figure 5) that allows a view on the results, helping singling out outliers and/or time intervals where the radio ephemeris was inadequate. Each data point shows the delay of the X-ray peak with respect to the radio peak for a single observation. Data from a given instrument can be shown or hidden by clicking on the legend. By hovering over each point, the user can see the observation ID, the mission, the TOA, the error on the TOA, the energy range, and a diagnostic plot showing if the fit was successful. Time intervals where the radio solution was inadequate (e.g. due to glitches) or absent (e.g., April 2020 due to the COVID pandemic) can be easily identified by the presence of a relatively large number of outliers from different missions.
4 Results
@@
Mission Instrument NN $r_{\rm mean}$ $\sigma$ $\sigma_{\rm stat}$
$\rm\mu s$ $\rm\mu s$ $\rm\mu s$
Chandra HRC 11 -503 350 25
Einstein Probe FXT 7 -391 79 15
Fermi LAT 203 -165 83 14
HITOMI HXI1 2 -264 27
HITOMI HXI2 2 -214 63
HITOMI SGD1 1 -244 78
HITOMI SGD2 1 -157 690
HITOMI SXS 1 -106 16
HXMT HE 343 -370 90 12
HXMT LE 386 -320a 100 62
HXMT ME 520 -333 85 18
IXPE DU1 9 -341 78 7.6
IXPE DU2 9 -344 74 8.6
IXPE DU3 9 -337 77 7.6
MAXI GSC 1337 -314 210 120
NICER XTI 145 -313 86 7
NINJASAT GMC 15 -406 54 22
NUSTAR FPMA 100 -363 86 12
NUSTAR FPMB 100 -361 76 13
SUZAKU HXD 18 -326 170 310
SWIFT BAT 3 20.5 160 58
SWIFT XRT 43 -615 150 1200
XMM EPN 98 -355 92 15
(a) We applied a known offset of-864 $\rm\mu s$} to the measured
value, (Y. Tuo et al., 2022)
@@
We applied the method described above to thousands of observations of the Crab pulsar involving 14 missions, spanning 23 years of observations, from 2002 to 2025.
@@
This wealth of data allows us to measure the relative timing accuracy of different missions, and to compare the results with the JBE (Figure 6). The X-ray TOAs are typically leading the radio TOAs, as reported in the past (e.g. A. H. Rots et al., 2004; D. A. Smith et al., 2008; G. Cusumano et al., 2012). The time difference between the X-ray and radio TOAs is not consistent with being constant, but it varies with the energy range of the photons, and notably, with time.
@@
We investigated the possibility that the method described in Section 3 introduces systematic errors in the TOAs. We verified that the scatter of the TOAs after the roundtrip procedure is typically of the order of the uncertainties of the simulated data (Figure 3). Moreover, we repeated the procedure using data barycentered with different JPL ephemerides and source positions, including the exact JBE position, and we found that the TOAs are consistent within 5 $\mu$}s or the uncertainties of the TOA calculation (Figure 4).
@@ 5 Conclusions
In this work, we presented a method to compare pulsar times of arrival (TOAs) from different space missions, even when using heterogeneous timing solutions, JPL ephemerides, and source positions. The method is based on the simulation of geocentered TOAs using a reference timing solution, and then fitting a new timing solution to these TOAs using the ephemerides and source position used for barycentering the data. We validated the method by simulating and comparing TOAs across different ephemerides (DE200–DE440), and verified that the TOAs are consistent within 5 $\rm\mu s$or the uncertainties of the TOA calculation. We applied the method to thousands of Crab pulsar observations from 14 missions over 23 years (2002–2025), enabling robust cross-calibration of timing accuracy. We find good agreement between the TOAs from GPS-carrying missions like NICER/XTI and HXMT/ME, with mean offsets within a few microseconds and scatter consistent with the expected statistical and systematic uncertainties. This demonstrates the robustness of the method for high-precision timing cross-calibration across modern missions. We released open-source tools and a TOA database to support future calibration and scientific studies using the Crab pulsar, and plan to extend the software and database to include additional missions and pulsars.
@@
MB wishes to acknowledge partial support by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). This work was supported by the JSPS Core-to-Core Program (grant number: JPJSCCA20220002) and Japan Society for the Promotion of Science Grants-in-Aid for Scientific Research (KAKENHI) Grant Number JP20K04009 (YT), JP19K14762, JP23K03459, JP24H01812 (MS). This work uses data obtained with Einstein Probe, a space mission supported by the Strategic Priority Program on Space Science of the Chinese Academy of Sciences, in collaboration with ESA, MPE, and CNES (Grant No. XDA15310303, No. XDA15310103, No. XDA15052100). VLK and AR acknowledge support from NASA contract to the Chandra X-ray Center NAS8-03060.
Appendix A Mission-specific data preparation
@@ A.1 Chandra/HRC
All the data are processed using the CIAO analysis software (v4.17 A. Fruscione et al., 2006) to apply the latest calibration, and the times are corrected for barycentric offsets using the CIAO tool axbary, using the JPL DE405 solar system ephemeris. Chandra’s spatial resolution is sufficient to isolate the pulsar as a point-like source within the diffuse nebula. We have estimated the position of the pulsar by centroiding on this point-like source and collect source photons within a circle of radius $1.5$ arcsec.
A.2 Einstein Probe/FXT
EP/FXT timing mode observations of Crab are summarized in Table,3. The observational data were processed using the FXT Data Analysis Software (FXTDAS) version 1.20, alongside the FXT calibration database (CALDB) version 1.20. The data reduction pipeline, denoted as fxtchain, was employed with precise source localization parameters: Right Ascension (R.A.) = 05h34m31.972s and Declination (Dec) = $22^{\circ}00^{{}^{\prime}}52.07^{{}^{\prime\prime}}$. The standard Good Time Interval (GTI) selection criteria, “ELV>5&&COR>6&&SAA==0&&DYE_ELV>30”, were applied. The events within the energy range of 0.5–10 keV were selected. The fxtbary tool was utilized to convert the time of each event to barycentric corrected time using ephemeris “de430_plus_MarsPC.bsp”.
@@ A.3 Fermi
A.4 Hitomi
We retrieved the final archival Crab data products for Hitomi from HEASARC. For the Soft X-ray Spectrometer (SXS), we utilized the cleaned event FITS file to extract the high- and medium-resolution grades (Hp, Mp, and Ms; Y. Ishisaki et al. (2018)) by selecting ”ITYPE=0:2” within xselect. The Hard X-ray Imager (HXI) data was obtained from the standard cleaned event FITS file, and extracted using a sky region around the Crab pulsar and extending up to a radius of 70 arcseconds from the image centroid. Only one unit of the Soft Gamma-ray Detector (SGD) is available, because the observation taking place during the commissioning phase. To increase the photon statistics, we selected photo-absorption event from the unscreened event FITS file, following the guidelines in Appendix 2 of Hitomi Collaboration (2018).
Subsequently, barycentric correction was applied on the DE200 solar system ephemeris for the event FITS files using the command outlined in Y. Terada et al. (2018):
@@ A.6 IXPE
A.7 MAXI/GSC
The MAXI/Gas Slit Camera (GSC) has been monitoring the Crab nebula continuously for over 16 years, since August 2009. We retrieved the GSC data using the FTOOL mxdownload_wget and generated daily cleaned event files with mxproduct, selecting events within a $1.6^{\circ}$ radius around the Crab. We then extracted events in the 2–20 keV energy range using xselect. Barycentric correction was applied using the FTOOL barycen, based on the DE-200 solar system ephemeris and the source position of (RA, Dec) = (83.633083, 22.014500). To ensure sufficient statistics, we excluded days with fewer than 1000 counts. As a result, we obtained event files for 4089 days, up to July 10, 2025.
A.8 NICER

### v3

@@ III.6 Pipeline
All steps above are automated in the TOAExtractor code [^11] https://matteobachetti.github.io/TOAextractor/ v. 0.6.0 (Bachetti, 2025c). This code is structured as a pipeline, using the luigi orchestrator [^12] https://luigi.readthedocs.io/en/stable/ v. 3.6.0. This software allows for a simple and modular way to define the steps of the pipeline, and to run them in parallel over multiple files by using multiple “workers”. The end of each step of the pipeline is marked by the production of a file. The appearance of this file triggers the start of the next step. Failing steps do not produce the output file, blocking the pipeline on a given file. Fixing the error and re-running the pipeline will restart the processing from the failed step. To reprocess a given dataset from one particular step of the pipeline, it is sufficient to delete all files produced by the step to be repeated and all subsequent steps, and re-run the pipeline.
III.7 Plotting and quality assurance
@@
Figure 5: Interactive summary plot produced by TOAExtractor. This plot allows for a simple exploration of the residuals, by showing diagnostic information as the user hovers over each point. The horizontal grey band indicates the X-ray residual measured by Rots et al. (2004). Vertical red lines indicate glitch epochs (In this static image it might be difficult to distinguish them from some large error bars from the data points, but the difference becomes clear in the interactive interface). Problematic observations and time intervals are easily identified: some observations with large residuals might be due to low count rates, temporary instrumental issues, or local issues with the radio ephemerides (e.g. the large deviation around April 2020, MJDs 58925–58975, due to the closure of the observatory for the COVID pandemic; recent glitches can also make the solution less reliable).
@@
This pipeline is conceived to be run on a large number of observations from different missions, possibly ever increasing over time. As such, it is important to have a quick way to check the quality of the data. For this reason, we developed an interactive summary plot (Figure 5) that allows a view on the results, helping to single out outliers and/or time intervals where the radio ephemeris was inadequate. Each data point shows the delay of the X-ray peak with respect to the radio peak for a single observation. Data from a given instrument can be shown or hidden by clicking on the legend. By hovering over each point, the user can see the observation ID, the mission, the TOA, the error on the TOA, the energy range, and a diagnostic plot showing if the fit was successful. Time intervals where the radio solution was inadequate (e.g. due to glitches) or absent (e.g., April 2020 due to the COVID pandemic) can be easily identified by the presence of a relatively large number of outliers from different missions.
IV Results
@@
Mission Instrument Min MJD Max MJD NN $r_{\rm mean}$ $\sigma$ $\sigma_{\rm stat}$ Rate Pulsed fraction
$\rm\mu s$ $\rm\mu s$ $\rm\mu s$ Counts s-1 %
AXAF/Chandra/CXO HRC 51574.28 56589.14 11 -458 310 83 2.42 100
EP FXT 60368.38 60758.73 7 -391 79 15 1490 34
GLAST/Fermi LAT 54696.08 60617.92 196 -167 80 14 0.00498 84
HITOMI HXI1 57472.63 57472.75 2 -264 27 73.1 61
HITOMI HXI2 57472.63 57472.74 2 -214 63 68.6 57
HITOMI SGD1 57472.63 57472.63 1 -244 78 1.17 52
HITOMI SGD2 57472.63 57472.63 1 -157 690 0.217 50
HITOMI SXS 57472.64 57472.64 1 -106 16 126 46
HXMT HE 57992.43 60398.80 342 -370 90 12 1360 36
HXMT LE 57992.60 60398.96 385 -319a 100 61 1220 38
HXMT ME 57992.26 60398.95 518 -332 86 18 439 37
IXPE DU1 59632.22 60727.75 7 -350 81 7.6 2.81 77
IXPE DU2 59632.22 60727.75 7 -352 76 8.5 2.5 75
IXPE DU3 59632.22 60727.75 7 -345 80 7.7 2.45 76
MAXI GSC 55135.51 60866.50 3847 -312 200 120 0.42 49
NICER XTI 57970.82 60364.52 145 -314 90 7 10400 38
NINJASAT GMC 60365.09 60771.04 14 -406 53 22 26 44
NUSTAR FPMA 56132.26 60388.08 79 -328 79 12 417 50
NUSTAR FPMB 56132.26 60388.08 78 -329 76 14 394 49
SUZAKU HXD 53604.23 56722.55 18 -326 170 310 53 39
SWIFT BAT 57431.04 58190.06 3 20.5 160 58 6770 6.7
SWIFT XRT 53454.21 58725.60 43 -615 150 1200 644 21
XMM EPN 51632.96 59269.92 97 -356 90 15 178 42
XRISM RESOLVE 60388.50 60388.50 1 -312 15 38.3 36
XTE PCA 50205.36 55927.00 312 -333 74 6.1 10600 46
(a) We applied a known offset of -864 $\rm\mu s$} to the measured value (Tuo et al., 2022)
@@
We applied the method described above to thousands of observations of the Crab pulsar involving 15 missions, spanning 29 years of observations, from 1996 to 2025.
@@
This wealth of data allows us to measure the relative timing accuracy of different missions, and to compare the results with the JBE (Figure 6). The X-ray TOAs are typically leading the radio TOAs, as reported in the past (e.g. Rots et al., 2004; Smith et al., 2008; Cusumano et al., 2012). The time difference between the X-ray and radio TOAs is not consistent with being constant, but it varies with the energy range of the photons, and notably, with time.
@@
We investigated the possibility that the method described in Section III introduces systematic errors in the TOAs. We verified that the scatter of the TOAs after the roundtrip procedure is typically of the order of the uncertainties of the simulated data (Figure 3). Moreover, we repeated the procedure using data barycentered with different JPL ephemerides and source positions, including the exact JBE position, and we found that the TOAs are consistent within 5 $\mu$}s or the uncertainties of the TOA calculation (Figure 4).
@@ V Conclusions
In this work, we presented a method to compare pulsar times of arrival (TOAs) from different space missions, even when using heterogeneous timing solutions, JPL ephemerides, and source positions. The method is based on the simulation of geocentered TOAs using a reference timing solution, and then fitting a new timing solution to these TOAs using the ephemerides and source position used for barycentering the data. We validated the method by simulating and comparing TOAs across different ephemerides (DE200–DE440), and verified that the TOAs are consistent within 5 $\rm\mu s$ or the uncertainties of the TOA calculation. We applied the method to thousands of Crab pulsar observations from 15 missions over 29 years (1996–2025), enabling robust cross-calibration of timing accuracy. We find good agreement between the TOAs from GPS-carrying missions like NICER/XTI and HXMT/ME, with mean offsets within a few microseconds relative to each other, and scatter consistent with the expected statistical and systematic uncertainties. This demonstrates the robustness of the method for high-precision timing cross-calibration across modern missions. We released open-source tools and a TOA database to support future calibration and scientific studies using the Crab pulsar, and plan to extend the software and database to include additional missions and pulsars.
@@
The authors wish to thank the anonymous referee for their insightful and constructive comments, which contributed greatly to the clarity and soundness of the paper. Thanks also to Andrew Lyne, Paul Ray, George Younes, David Smith, Allyn Tennant for thoughtful discussions. MB wishes to acknowledge partial support by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). This work was supported by the JSPS Core-to-Core Program (grant number: JPJSCCA20220002) and Japan Society for the Promotion of Science Grants-in-Aid for Scientific Research (KAKENHI) Grant Number JP20K04009 (YT), JP19K14762, JP23K03459, JP24H01812 (MS). This work uses data obtained with Einstein Probe, a space mission supported by the Strategic Priority Program on Space Science of the Chinese Academy of Sciences, in collaboration with ESA, MPE, and CNES (Grant No. XDA15310303, No. XDA15310103, No. XDA15052100). VLK and AR acknowledge support from NASA contract to the Chandra X-ray Center NAS8-03060. This work made significant use of the Fornax Science Console, which is part of the NASA Astrophysics cloud-based Fornax Initiative jointly developed by Goddard Space Flight Center’s Astrophysics Projects Division (GSFC/ApPD) and the astrophysics archives – the High Energy Astrophysics Science Archive Research Center (HEASARC), the Infrared Science Archive (IRSA), and the Mikulski Archive for Space Telescopes (MAST).
Appendix A Time definitions and conventions
Figure 7: Relevant time scales and transformations described in Section A.
When doing precise timing work with pulsars, it is important to agree on two fundamental aspects: the time scale (which decides how fast our clocks run and a specific reference time to refer our time measurements to), and the relative position of the observer, the pulsar, and the solar system barycenter (which allows us to correct for different light travel times from the source to our instrument as the Earth moves around its orbit). What we summarize here is treated in great detail in the TEMPO2 paper II by Edwards et al. (2006). There is also some confusion on the acronyms used by technical reports on time scales and in pulsar astronomy papers, that we try to address here.
A.1 Time scales
The calendar utilized in everyday tasks is based on the duration of a day, which is defined as a rotation of the Earth with respect to the Sun131313not a full rotation on its axis, which would be a sidereal day, and initially second was defined as 1/86400th of the duration of an average day. However, this process does not last a constant amount of time, but the Earth’s rotation is not stable and, over time, tends to slows down, mostly due to the tidal effect of the Moon. Since 1967, the second has become an International System of Units (SI) base unit, defined as the duration of 9,192,631,770 periods of the radiation corresponding to the transition between the two hyperfine levels of the ground state of the 133Cs atom (Bureau International des Poids et Mesures, 2025). A day is now defined as 86400 s, and to correct for the Earth rotation evolution, leap seconds are introduced from time to time to synchronize the calendar to the rotation (so that there are single days of 86401 seconds, like December 31, 2016). The time scale tied to the Earth rotation is called Coordinated Universal Time (UTC). It is calculated as a byproduct of the Terrestrial Time (TT), which is a theoretical time measured on the surface of a spherical Earth whose radius is the mean sea level. The most commonly used estimate (or realization) of the TT scale is the international atomic time (TAI), calculated from the weighted (and corrected for relativistic effects) average of 450 atomic clocks, mostly caesium clocks. The Bureau International de Poids et Measures (BIPM) annually reanalizes the same clock time series used by TAI and produces an improved TT solution that can be used for the most challenging metrology applications. Independent realization of TT are under study, notably one based on pulsars (Hobbs et al., 2020), which would definitively disentangle our measurement of time from the Earth’s influence and be suited to interstellar travel.
Because of general relativity, the pace at which clocks run is a function of their acceleration (or, equivalently, the gravitational field they are subject to). A clock at rest in empty space will run faster than one close, e.g., to Earth; different clocks at different positions on the geoid will run at different paces, being subject to a slightly different gravitational pull; and as the Earth moves around the Sun in its elliptical orbit, the rate at which the ideal clock from TT runs remains non constant, due to the distance from the Sun. This effect is called Einstein delay (Damour and Deruelle, 1986). This gives rise to a periodic change of clock rates which needs to be taken into account in precision timing. The time scale correcting TT for these relativistic rate changes is called barycentric dynamical time (TDB). An even more fundamental scale is the barycentric coordinate time (TCB), which is equivalent to the proper time measured by a clock which is co-moving with the barycenter of the Solar system but outside its gravitational well, calculated in a fully relativistic framework. In practice, TCB differs from TDB by a fixed rate [^14] IAU 2006 Resolution 3: https://syrte.obspm.fr/iauJD16/IAU2006_Resol3.pdf, and TDB is nowadays defined in terms of TCB. The Tempo2 package (Hobbs et al., 2006) internally uses TCB, while TEMPO (Nice et al., 2015) and PINT (Luo et al., 2021) use TDB. TDB is also still used as an independent variable in the JPL-DE ephemerides. In parallel, the Solar System Einstein delay is computed using the IF99 model (Irwin and Fukushima, 1999) by Tempo2 and the FB90 (Fairhead and Bretagnon, 1990) by TEMPO and PINT [^15] They are specified through the TIMEEPH parameter in parameter files.
As a side note, despite the name similarity, TDB is a time scale valid everywhere in the relativistic framework. It does not necessarily involve the so-called barycentering of the data, which consists of correcting the time measurements for light travel time across the solar system (see below).
A.2 Observer position-related transformations
Once we have agreed on a time scale and our clocks are synchronized, we still need to deal with the fact that signals travel in space at a finite velocity, the speed of light.
Depending on the position of the Earth in its orbit, the arrival time at the instrument of a given photon can change by up to 16 minutes (Roemer delay). The same applies to smaller but significant corrections by milliseconds depending on the instrument position on the Earth surface or in orbit. It is therefore important to find a way to express the time of arrival at a fixed reference point, typically the solar system barycenter (SSB). Correcting for this effect is called barycentering.
Modern pulsar timing codes use the detailed models of the solar system provided by the JPL-DE in order to determine not only the position of the SSB, but also other kinds of delay due to the signal traversing the curved space time near the Sun and other Solar System objects, the so-called Shapiro Delay (Shapiro, 1964; Damour and Deruelle, 1986).
Appendix B Mission-specific data preparation
@@ B.1 Chandra/HRC
All the data are processed using the CIAO analysis software (v4.17 Fruscione et al., 2006) to apply the latest calibration, and the times are corrected for barycentric offsets using the CIAO tool axbary, using the JPL-DE405 solar system ephemeris. Chandra’s spatial resolution is sufficient to isolate the pulsar as a point-like source within the diffuse nebula. We have estimated the position of the pulsar by centroiding on this point-like source and collect source photons within a circle of radius $1.5$ arcsec.
B.2 Einstein Probe/FXT
EP/FXT timing mode observations of Crab are summarized in Table,4. The observational data were processed using the FXT Data Analysis Software (FXTDAS) version 1.20, alongside the FXT calibration database (CALDB) version 1.20. The data reduction pipeline, denoted as fxtchain, was employed with precise source localization parameters: Right Ascension (R.A.) = 05h34m31.972s and Declination (Dec) = $22^{\circ}00^{{}^{\prime}}52.07^{{}^{\prime\prime}}$. The standard Good Time Interval (GTI) selection criteria, “ELV>5&&COR>6&&SAA==0&&DYE_ELV>30”, were applied. The events within the energy range of 0.5–10 keV were selected. The fxtbary tool was utilized to convert the time of each event to barycentric corrected time using ephemeris “de430_plus_MarsPC.bsp”.
@@ B.3 Fermi
B.4 Hitomi
We retrieved the final archival Crab data products for Hitomi from HEASARC. For the Soft X-ray Spectrometer (SXS), we utilized the cleaned event FITS file to extract the high- and medium-resolution grades (Hp, Mp, and Ms; Ishisaki et al. (2018)) by selecting "ITYPE=0:2" within xselect. The Hard X-ray Imager (HXI) data was obtained from the standard cleaned event FITS file, and extracted using a sky region around the Crab pulsar and extending up to a radius of 70 arcseconds from the image centroid. Only one unit of the Soft Gamma-ray Detector (SGD) is available, because the observation taking place during the commissioning phase. To increase the photon statistics, we selected photo-absorption event from the unscreened event FITS file, following the guidelines in Appendix 2 of Hitomi Collaboration (2018).
Subsequently, barycentric correction was applied on the DE200 solar system ephemeris for the event FITS files using the command outlined in Terada et al. (2018):
@@ B.6 IXPE
B.7 MAXI/GSC
The MAXI/Gas Slit Camera (GSC) has been monitoring the Crab nebula continuously for over 16 years, since August 2009 (Morii et al., 2011; Sugizaki et al., 2011). We retrieved the GSC data using the FTOOL mxdownload_wget and generated daily cleaned event files with mxproduct, selecting events within a $1.6^{\circ}$ radius around the Crab [^18] For the details of MAXI analysis, see https://darts.isas.jaxa.jp/missions/maxi/analysis/. We then extracted events in the 2–20 keV energy range using xselect. Barycentric correction was applied using the FTOOL barycen, based on the DE-200 solar system ephemeris and the source position of (RA, Dec) = (83.633083, 22.014500). To ensure sufficient statistics, we excluded days with fewer than 1000 counts. As a result, we obtained event files for 4089 days, up to July 10, 2025.
B.8 NICER